Darla is building a new desk. To make sure she had made a square corner, she measures 4 ft. from the corner along one edge and 9 ft. from the corner along the other edge. How long should the diagonal be between those two points if the corner is a right angle? Diagonal = how many feet?
Answer by stanbon(75887)

Darla is building a new desk. To make sure she had made a square corner, she measures 4 ft. from the corner along one edge and 9 ft. from the corner along the other edge. How long should the diagonal be between those two points if the corner is a right angle? Diagonal = how many feet?
---------------
Use Pythagoras:
d^2 = 4^2 + 9^2
d^2 = 36+81
d^2 = 117
d = 10.82 ft
===================
Cheers,
Stan H.